Answered

Two students from a group of eight boys and 12 girls are sent to represent the school in a parade.


If the students are chosen at random, what is the probability that the students chosen are not both girls?

A12/190
B33/95
C62/95
D178/190

Verified Answer :

Answer

c. 62/95


Explanation

Probability is the chance of an event happening.

Probability = (favorable outcome)/(total outcome)

P(girl) = 12/20 × 11/19

        = (12×11)/(20×19)

         = 132/380

         = 33/95

P( not a girl) = 1 - 33/95

                     = 62/95

Answer:

62/95

Step-by-step explanation:

none